0.23 kg of air has an initial pressure of 1.7 MN/m2 and a temperature of 2000C. It is expanded to a pressure of 0.34MN/m2 according to the law PV1.35 = constant. determine the work transfer during the expansion. take R= 0.29 kj/kg K


1
Expert's answer
2021-10-14T23:42:01-0400

For any polytropic process;

"PV^{n}=C"

"W=\frac{P_1V_1-P_2V_2}{n-1}=\frac{mR(T_1-T_2)}{n-1}"

We first find T2;

"PV^{1.35}=C"

"V \propto\frac{T}{P}"

"P(\frac{T}{P})^{1.35}=C"

"P^{1-1.35}T^{1.35}=C"

"T^{1.35}\propto P^{0.35}"

"(\frac{T_2}{T_1})^{1.35}=(\frac{P_2}{P_1})^{0.35}"

"T_2=T_1(\frac{P_2}{P_1})^{\frac{0.35}{1.35}}"

"=473(\frac{0.34}{0.17})^{\frac{0.35}{1.35}}"

=566.116K

"W=\frac{mR(T_1-T_2)}{n-1}"

"W=\frac{0.23*0.29(566.116-473)}{1.35-1}"

"W=12.745kJ"